Doggy Treats is selling dog terats in a purely competitive market. Its output is 800 treats, which it sells for $10 a treat. At the 800-treat leel of output, the marginal cost is $11, the average variable cost is $9.00, and the average variable is $8.00. Should the firm increase output, decrease output, or not produce? Why? How should the firm determine that optimal level of output.
1
Expert's answer
2012-11-21T08:33:49-0500
TP = (P - ATC)*Q = (P - AVC - AFC)*Q = (10 - 9 -8)*800 = -$5600
The firm faces losses, but its ATC<P<AVC, so it should work.
It should decrease its production as P = MR < MC